high temp in a certain city for the month of August follow a unifrom distribution over the interval 65 degree F to 88 degree F. What is the probability that a randomly selected August day has a high temp that exceeded 70 degree F?
0.4575
0.0435
0.7826
0.2174
Answer #1

Solution:

X follows Uniform(a = 65 , b = 88)

PDF of X is

f(x) = 1/(b - a) = 1/(88 - 65) = 1/23

Now ,

P(exceeded 70)

= P(X > 70)

= (b - 70)/(b - a)

= (88 - 70)/(88 - 65)

= 18/23

= 0.7826